OpenStudy (anonymous):

f(x) = x 2 + 1 for f(-1).

OpenStudy (austinl):

\(f(x)=x^2+1\) For f(-1) just input -1 for all instances of x in the equation.

OpenStudy (anonymous):

is that it?

OpenStudy (austinl):

@Haseeb96 They made an error in formatting. it is x^2

OpenStudy (haseeb96):

f(x)=x^2+1 For f(-1) f(-1)=(-1)^2+1 f(-1)=+1+1 f(-1)=+2 this is the solution and it is absolutely correct